# Benjamin Carleton

### v6

@@ 2.2 Control Types and Collapses
Here are the collapses found in Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [${}^{+}$22]. For every election system $\mathcal{E}$, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ and -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ ([HHM20] supplemented by [${}^{+}$22]’s observation that the [HHM20] collapse proofs work not just over linear orders but in fact work over any election system regardless of its vote type). Property Unique-$\alpha$ [HHR07], the unique version of Chernoff’s Property $\alpha$ ([Che54], see [Sen71]), is said to hold for election system $\mathcal{E}$ if in every election $(C,V)$, a candidate $p$ being a unique winner implies that $p$ remains a unique winner whenever the election is restricted to a subset $C^{\prime}\subseteq C$ of candidates such that $p\in C^{\prime}$. For every election system $\mathcal{E}$ satisfying Property Unique-$\alpha$, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ and -$\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ [${}^{+}$22]. [^4] What we call Property Unique-$\alpha$ (and for brevity we often drop the “Property”) has been previously referred to as the “Unique-WARP” property. We discuss below our reasons for moving away from that name. Bartholdi et al. [BTT92] first used the term WARP (Weak Axiom of Revealed Preferences) in the domain of computational social choice by stating that it required that $p$ being a winner of election $(C,V)$ implies that $p$ remains a winner of every election $(C^{\prime},V)$, where $p\in C^{\prime}\subseteq C$, and they stated that “this” was also known as Property $\alpha$. Unfortunately, Hemaspaandra et al. [HHR07] apparently read that as saying that a definition of WARP was being given. And so they used the term “Unique-WARP” when they defined the variant of that in which $p$ is required to be a unique winner. In fact, WARP and Property $\alpha$ are not equivalent. (Sen [Sen71], however, proved that WARP—in its long-settled and standard sense, see [HG22, Section 5.1]—is, with respect to the standard notion of choice functions (election systems) used in pure social choice theory, equivalent to the combination of Property $\alpha$ and a so-called Property $\beta$ (from [Sen69]) that we will not define here. The reason we mention that the equivalence holds for the pure social choice notion of election systems is that in our model, where empty winner sets are allowed, the equivalence in fact fails. Consider the following election system $\mathcal{E}$. If three or more candidates run, then the two candidates with the lexicographically smallest names win. Otherwise, no one wins. Clearly, this election system does not satisfy Property $\alpha$ (and certainly not Properties $\alpha$ and $\beta$ combined), but it does satisfy WARP.) To summarize briefly, a misreading by [HHR07] of an ambiguous sentence in [BTT92] led to an infelicitous naming in [HHR07] that we feel should be abandoned. We thus use “Unique-$\alpha$” to denote what to date has been called “Unique-WARP” (e.g., in [HHR07, Erd09, ENR09, EFRS15, FHH20, ${}^{+}$22]). We do that so the terminology is analogous to what has long been used in social choice for the notion that inspired the Unique-$\alpha$ notion. Since approval voting satisfies Unique-WARP, these four additional pair-collapses hold for approval voting [${}^{+}$22]. Note that $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ participates in both the four-type and the three-type above collapses with $\mathcal{E}={\rm Approval}$, so by transitivity we have the six-type collapse $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ [${}^{+}$22].
@@
The remaining election-specific collapses from Carleton et al. [${}^{+}$22] are $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, and $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$.
@@ 2.3 Search Problems and Their Interreductions and Complexity
For each election system $\mathcal{E}$, define the winner problem, $W_{\mathcal{E}}$, by $W_{\mathcal{E}}=\{(C^{\prime},V^{\prime},p^{\prime})\mid p^{\prime}\in C^{\prime}$ and $p^{\prime}$ is a winner of the $\mathcal{E}$ election $(C^{\prime},V^{\prime})\}$. Our control types are each defined as a language problem—a set. However, each has a clear “$\mathrm{NP}^{W_{\mathcal{E}}}$ search problem” associated with it (so an “$\mathrm{NP}$ search problem” if $\mathcal{E}$’s winner problem is in $\mathrm{P}$). Namely, the problem of, on input $(C,V,p)$, outputting a partition $(C_{1},C_{2})$ of the candidate set (or if the type is a voter partition type, a partition $(V_{1},V_{2})$ of the vote collection) such that under that partition, the control action regarding $(C,V,p)$ is successful under election system $\mathcal{E}$.
To be able to speak clearly of search problems and their interrelationships and complexity it is important to be specific as to what we mean, both in terms of relating the solutions of two search problems and as to classifying the complexity of a search problem. These two different tasks are related, yet differ in how they are formalized.
@@
In the case of our 24 electoral-control problems, the “$x$” here is the problem input, $(C,V,p)$. And if $(C,V,p)$ is a yes instance of the (viewed as a set) control type, an acceptable output $y$ in the sense of the above definition is (the encoding of) any partition $(C_{1},C_{2})$ of $C$ (or if the type is a voter partition, then any partition $(V_{1},V_{2})$ of $V$) that leads to success under the control type, the input $(C,V,p)$, and the election system. (We in this paper will not worry about encoding details, since such details are not a key issue here. We as is typical merely assume reasonable, natural encodings.)
@@
In what follows, we will reformulate the definition more carefully, for our case, to close that loophole (which is easily closed). However, more interestingly, we will in effect alter the definition in two additional ways. First, notice that in Megiddo and Papadimitriou’s definition (assuming one fixes the above loophole first) the function $f(x)$ allows the solution to $x$ on the $R$ side to be obtained via demanding a solution to a different (than $x$) instance $f(x)$, on the $S$ side. But in our setting, collapsing types are the same set, just with differing “witnessing” relations. And our goal is to make connections via those witnesses. So in the definitions we are about to give of our “$\leq_{\textnormal{search}}$” reduction family, we in effect require their $f(x)$ to be the identity function! Second, since we wish to connect the solutions of collapsing pairs even when the winner problem of $\mathcal{E}$ is not in $\mathrm{P}$, our reductions will have the winner problem as an oracle. (The reason for this is that even given a $(C,V,p)$ and a partition $(C_{1},C_{2})$ of $C$ (or for voter partition types, a partition $(V_{1},V_{2})$ of $V$), to evaluate whether the partition is a solution generally needs calls to $W_{\mathcal{E}}$.)
We will not in these definitions explicitly write “$\Pi$” expressions, but the solutions we speak of are with respect to the witnesses/actions that the given control type is about (they are the search component of $\Pi$, i.e., in our case they are those second components that appear in pairs, belonging to the underlying relation $R_{\mathcal{T}}$, having “$x$” or “$I$” (i.e., the given $(C,V,p)$) as their first component), and so we view these definitions as a reformulated version (with some changes for our particular situation) of the Megiddo and Papadimitriou “$\Pi$” approach.
@@
    For an election system $\mathcal{E}$ and $\mathcal{E}$-matched, collapsing [^5] The reason we include “collapsing” in the definition is that if it is omitted, then one would trivially satisfy the notion on a given $I$ whenever $I$ was not in $\mathcal{T}_{2}$. That is, our notion is focused on pairs of types that collapse—where each input is either in both or not in both. control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we say that “$\mathcal{T}_{1}$ is polynomially search-reducible to $\mathcal{T}_{2}$ with respect to $\mathcal{E}$” (denoted by $\mathcal{T}_{1}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{2}$) if there is a reduction that, given an oracle for the winner problem for $\mathcal{E}$, runs in polynomial time and on each input $(I,S)$, where $I$ is an input to $\mathcal{T}_{1}$ and $S$ is a solution for $I$ with respect to $\mathcal{T}_{2}$, outputs a solution $S^{\prime}$ for $I$ with respect to $\mathcal{T}_{1}$.
@@
    For an election system $\mathcal{E}$ and $\mathcal{E}$-matched, collapsing control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we say that “$\mathcal{T}_{1}$ is polynomially search-equivalent to $\mathcal{T}_{2}$ with respect to $\mathcal{E}$” if $\mathcal{T}_{1}$ is polynomially search-reducible to $\mathcal{T}_{2}$ with respect to $\mathcal{E}$ and $\mathcal{T}_{2}$ is polynomially search-reducible to $\mathcal{T}_{1}$ with respect to $\mathcal{E}$.
@@
    For an election system $\mathcal{E}$ and $\mathcal{E}$-matched, collapsing control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we say that “$\mathcal{T}_{1}$ is polynomially search-reducible to $\mathcal{T}_{2}$” (denoted by $\mathcal{T}_{1}\leq_{\textnormal{search}}^{p}\mathcal{T}_{2}$) if there is a reduction that runs in polynomial time and on each input $(I,S)$, where $I$ is an input to $\mathcal{T}_{1}$ and $S$ is a solution for $I$ with respect to $\mathcal{T}_{2}$, outputs a solution $S^{\prime}$ for $I$ with respect to $\mathcal{T}_{1}$.
@@
    For an election system $\mathcal{E}$ and $\mathcal{E}$-matched, collapsing control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we say that “$\mathcal{T}_{1}$ is polynomially search-equivalent to $\mathcal{T}_{2}$” if $\mathcal{T}_{1}$ is polynomially search-reducible to $\mathcal{T}_{2}$ and $\mathcal{T}_{2}$ is polynomially search-reducible to $\mathcal{T}_{1}$.

### v7

@@ 2.2 Control Types and Collapses
Before giving the collapses found by Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [CCH+24], we give the definitions of three axiomatic properties of election systems. An election system $\mathcal{E}$ is said to satisfy Property $\alpha$ (see [Che54, Sen71]) if for every election $(C,V)$ and every candidate $p\in C$ it holds that if $p$ is a winner of election $(C,V)$ then for every $C^{\prime}$ satisfying $p\in C^{\prime}\subseteq C$ candidate $p$ remains a winner in the restriction of $(C,V)$ to $C^{\prime}$. An election system $\mathcal{E}$ is said to satisfy Property Unique-$\alpha$, the unique version of Property $\alpha$, if for every election $(C,V)$ and every candidate $p\in C$ it holds that if $p$ uniquely wins election $(C,V)$ then for every $C^{\prime}$ satisfying $p\in C^{\prime}\subseteq C$ candidate $p$ uniquely wins in the restriction of $(C,V)$ to $C^{\prime}$. [^5] What we call Property Unique-$\alpha$ (and for brevity we often drop the “Property”) has been previously referred to as the “Unique-WARP” property. We discuss below our reasons for moving away from that name. Bartholdi et al. [BTT92] first used the term WARP (Weak Axiom of Revealed Preferences) in the domain of computational social choice by stating that it required that $p$ being a winner of election $(C,V)$ implies that $p$ remains a winner of every election $(C^{\prime},V)$, where $p\in C^{\prime}\subseteq C$, and they stated that “this” was also known as Property $\alpha$. Unfortunately, Hemaspaandra et al. [HHR07a] apparently read that as saying that a definition of WARP was being given. And so they used the term “Unique-WARP” when they defined the variant of that in which $p$ is required to be a unique winner. In fact, WARP and Property $\alpha$ are not equivalent. (Sen [Sen71], however, proved that WARP—in its long-settled and standard sense, see [HG22, Section 5.1]—is, with respect to the standard notion of choice functions (election systems) used in pure social choice theory, equivalent to the combination of Property $\alpha$ and a so-called Property $\beta$ (from [Sen69]) that we will not define here. The reason we mention that the equivalence holds for the pure social choice notion of election systems is that in our model, where empty winner sets are allowed, the equivalence in fact fails. Consider the following election system $\mathcal{E}$. If three or more candidates run, then the two candidates with the lexicographically smallest names win. Otherwise, no one wins. Clearly, this election system does not satisfy Property $\alpha$ (and certainly not Properties $\alpha$ and $\beta$ combined), but it does satisfy WARP.) To summarize briefly, a misreading in [HHR07a] of an ambiguous sentence in [BTT92] led to an infelicitous naming in [HHR07a] that we feel should be abandoned. We thus use “Unique-$\alpha$” to denote what to date has been called “Unique-WARP” (e.g., in [HHR07a, Erd09, ENR09, EFRS15, FHH20, CCH+23b]). We do that so the terminology is analogous to what has long been used in social choice for the notion that inspired the Unique-$\alpha$ notion. Finally, an election system $\mathcal{E}$ is said to be strongly voiced exactly if for every election $(C,V)$ with $C\neq\emptyset$, there is at least one winner of $(C,V)$ under $\mathcal{E}$ [HHR07b].
Here are the collapses found in Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [CCH+24]. For every election system $\mathcal{E}$, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ and -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ ([HHM20] supplemented by [CCH+24]’s observation that the [HHM20] collapse proofs work not just over linear orders but in fact work over any election system regardless of its vote type). For every election system $\mathcal{E}$ satisfying Property Unique-$\alpha$, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ [CCH+24]. And for every election $\mathcal{E}$ that is strongly voiced and satisfies Property $\alpha$, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ [CCH+24]. Since approval voting is clearly strongly voiced and clearly satisfies Property $\alpha$ and Property Unique-$\alpha$, these four additional pair-collapses hold for approval voting [CCH+24]. Note that $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ participates in both the four-type and the three-type above collapses with $\mathcal{E}={\rm Approval}$, so by transitivity we have the six-type collapse $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=$ $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ [CCH+24].
@@
The remaining collapses from Carleton et al. [CCH+24] are $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, and finally, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}=\allowbreak\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$.
@@ 2.3 Search Problems and Their Interreductions and Complexity
For each election system $\mathcal{E}$, define the winner problem, $W_{\mathcal{E}}$, by $W_{\mathcal{E}}=\{(C^{\prime},V^{\prime},p^{\prime})\mid p^{\prime}\in C^{\prime}$ and $p^{\prime}$ is a winner of the $\mathcal{E}$ election $(C^{\prime},V^{\prime})\}$. Our control types are each defined as a language problem—a set. However, each has a clear $\mathrm{NP}^{W_{\mathcal{E}}}$ search problem” associated with it (so an $\mathrm{NP}$ search problem” if $\mathcal{E}$’s winner problem is in $\mathrm{P}$). Namely, the problem of, on input $(C,V,p)$, outputting a partition $(C_{1},C_{2})$ of the candidate set (or if the type is a voter partition type, a partition $(V_{1},V_{2})$ of the vote collection) such that under that partition, the control action regarding $(C,V,p)$ is successful under election system $\mathcal{E}$.
To be able to speak clearly of search problems and their interrelationships and complexity, it is important to be specific as to what we mean, both in terms of relating the solutions of two search problems and as to classifying the complexity of a search problem. These two different tasks are related, yet differ in how they are formalized.
@@
In the case of our 24 electoral-control problems, the “xx” here is the problem input, $(C,V,p)$. And if $(C,V,p)$ is a yes instance of the (viewed as a set) control type, an acceptable output $y$ in the sense of the above definition is (the encoding of) any partition $(C_{1},C_{2})$ of $C$ (or if the type is a voter partition, then any partition $(V_{1},V_{2})$ of $V$) that leads to success under the control type, the input $(C,V,p)$, and the election system. (In this paper we will not worry about encoding details, since such details are not a key issue here. We, as is typical, merely assume reasonable, natural encodings.)
@@
In what follows, we will reformulate the definition more carefully, for our case, to close that loophole (which is easily closed). However, more interestingly, we will in effect alter the definition in two additional ways. First, notice that in Megiddo and Papadimitriou’s definition (assuming one fixes the above loophole first) the function $f(x)$ allows the solution to $x$ on the $R$ side to be obtained via demanding a solution to a different (than $x$) instance $f(x)$, on the $S$ side. But in our setting, collapsing types are the same set, just with differing “witnessing” relations. And our goal is to make connections via those witnesses. So in the definitions we are about to give of our $\leq_{\textnormal{search}}$” reduction family, we in effect require their $f(x)$ to be the identity function! Second, since we wish to connect the solutions of collapsing pairs even when the winner problem of $\mathcal{E}$ is not in $\mathrm{P}$, our reductions will have the winner problem as an oracle. (The reason for this is that even given a $(C,V,p)$ and a partition $(C_{1},C_{2})$ of $C$ (or for voter partition types, a partition $(V_{1},V_{2})$ of $V$), to evaluate whether the partition is a solution generally needs calls to $W_{\mathcal{E}}$.)
We will not in these definitions explicitly write $\Pi$” expressions, but the solutions we speak of are with respect to the witnesses/actions that the given control type is about (they are the search component of $\Pi$, i.e., in our case they are those second components that appear in pairs, belonging to the underlying relation $R_{\mathcal{T}}$, having “xx” or “II” (i.e., the given $(C,V,p)$) as their first component), and so we view these definitions as a reformulated version (with some changes for our particular situation) of the Megiddo and Papadimitriou $\Pi$” approach.
@@
    For an election system $\mathcal{E}$ and $\mathcal{E}$-matched, collapsing [^6] The reason we include “collapsing” in the definition is that if it is omitted, then one would trivially satisfy the notion on a given $I$ whenever $I$ was not in $\mathcal{T}_{2}$. That is, our notion is focused on pairs of types that collapse—where each input is either in both or in neither. control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we say that $\mathcal{T}_{1}$ is polynomially search-reducible to $\mathcal{T}_{2}$ with respect to $\mathcal{E}$” (denoted by $\mathcal{T}_{1}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{2}$) if there is a reduction that, given an oracle for the winner problem for $\mathcal{E}$, runs in polynomial time and on each input $(I,S)$, where $I$ is an input to $\mathcal{T}_{1}$ and $S$ is a solution for $I$ with respect to $\mathcal{T}_{2}$, outputs a solution $S^{\prime}$ for $I$ with respect to $\mathcal{T}_{1}$.
@@
    For an election system $\mathcal{E}$ and $\mathcal{E}$-matched, collapsing control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we say that $\mathcal{T}_{1}$ is polynomially search-equivalent to $\mathcal{T}_{2}$ with respect to $\mathcal{E}$” if $\mathcal{T}_{1}$ is polynomially search-reducible to $\mathcal{T}_{2}$ with respect to $\mathcal{E}$ and $\mathcal{T}_{2}$ is polynomially search-reducible to $\mathcal{T}_{1}$ with respect to $\mathcal{E}$.
@@
    For an election system $\mathcal{E}$ and $\mathcal{E}$-matched, collapsing control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we say that $\mathcal{T}_{1}$ is polynomially search-reducible to $\mathcal{T}_{2}$” (denoted by $\mathcal{T}_{1}\leq_{\textnormal{search}}^{p}\mathcal{T}_{2}$) if there is a reduction that runs in polynomial time and on each input $(I,S)$, where $I$ is an input to $\mathcal{T}_{1}$ and $S$ is a solution for $I$ with respect to $\mathcal{T}_{2}$, outputs a solution $S^{\prime}$ for $I$ with respect to $\mathcal{T}_{1}$.
@@
    For an election system $\mathcal{E}$ and $\mathcal{E}$-matched, collapsing control types $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we say that $\mathcal{T}_{1}$ is polynomially search-equivalent to $\mathcal{T}_{2}$” if $\mathcal{T}_{1}$ is polynomially search-reducible to $\mathcal{T}_{2}$ and $\mathcal{T}_{2}$ is polynomially search-reducible to $\mathcal{T}_{1}$.